Hello Tutor, I really need your help solving this question please.
Solve the system: y = x2 + 5
6x – y = 3

Found 2 solutions by Boreal, fcabanski:
Answer by Boreal(15235)

I am going to do this two ways, one assuming it is y=x^2 +5, the other assuming it is y=2x+5. Both of these actually can be solved. The second is first.
Substitute
6x-(2x+5)=3
6x-2x-5=3
4x-5=3
4x=8
x=2
y=9
Other equation
6(2)-9=12-9=3
If it really were y=x^2 +5
6x-(x^2+5)=3
6x-x^2-5=3
-x^2+6x-8=0
x^2-6x+8=0
(x-4)(x-2)=0\
x=4, 2 y=21 ,9
Second equation is 24-21=3
12-9=3
Solution set for that is (4,21) and (2,9)


Answer by fcabanski(1391)